Calculate the volume percentage change as iron transform from FCC to BCC if the atomic radii are 1.270A and 1.241 A respectively

1
Expert's answer
2022-01-20T19:34:01-0500

In BCC there are 2 atoms per unit cell,

"a = \frac{4}{\sqrt3}r =\frac{4}{\sqrt3}\times1.270 =2.93"

an FCC crystal structure, 4 atoms per unit cell and

"a = 2 \sqrt2r =2\sqrt2\times1.241 =3.51"


percentage change as iron transforms from FCC to BCC "=\frac{3.51-2.93}{100}= 58 %" %